2x + y = 8 y= x - 7
Answer by mananth(16949)

2 x + 1 y = 8 .............1
-1 x + 1 y = -7 .............2
Eliminate y
multiply (1)by -1
Multiply (2) by 1
-2 x -1 y = -8
-1 x + 1 y = -7
Add the two equations
-3 x = -15
/ -3
x = 5
plug value of x in (1)
2 x + 1 y = 8
10 + 1 y = 8
1 y = 8 -10
1 y = -2
y = -2
(-3,-2)